OpenStudy (anonymous):

Can any body show me how?? [integrating from 0 to t]{f(t)*g(t-@)} d@=[integrating from 0 to infnity] {f(t)*g(t-@)*u(t-@) }d@ Here u(t) is the unit step function

myininaya (myininaya):

this is hard to read

myininaya (myininaya):

\[\int\limits_{0}^{t}f(t)g(t-\theta) d \theta=\int\limits_{0}^{\infty}f(t) g(t-\theta) u(t-\theta) d \theta\]

myininaya (myininaya):

is that right?

OpenStudy (anonymous):

yes

myininaya (myininaya):

i never heard of the unit step function

myininaya (myininaya):

oh thats the heavside function?

OpenStudy (anonymous):

u(t-a)={1 for t>a 0 t<a ,a>=0

OpenStudy (anonymous):

yes it is

myininaya (myininaya):

so this is like partial different equations or differential equations class?

OpenStudy (anonymous):

this is not any class I see that step in one of my class and I'm trying to understand how it's done

OpenStudy (anonymous):

soory of my book

myininaya (myininaya):

what book?

myininaya (myininaya):

oh lol i don't have that book i'm still thinking ok...

myininaya (myininaya):

yeah i just seen alot about laplace

myininaya (myininaya):

lets see if this helps http://www.physicsforums.com/showthread.php?t=434008 i'm still reading this so it may not help us

OpenStudy (anonymous):

This'll not help if you see defination of unit step then you can easily answer his post

myininaya (myininaya):

you need zarkon let me see if hes online

OpenStudy (anonymous):

If we break into two intervals [integrating from 0 to @] f(t)*g(t-@)*0d@+[integrating from @ to infnity] f(t)*g(t-@)*1d@=[integrating from @ to infnity] f(t)*g(t-@)d@ I'm not getting how can be it equal to [integrating from 0 to t]{f(t)*g(t-@)} d@

OpenStudy (zarkon):

\[u(t−\theta)=1\] if \[t-\theta\ge 0\] and 0 otherwise \[\int\limits_{0}^{\infty}f(t) g(t-\theta) u(t-\theta) d \theta\] \[=\int\limits_{0}^{t}f(t) g(t-\theta) u(t-\theta) d \theta+\int\limits_{t}^{\infty}f(t) g(t-\theta) u(t-\theta) d \theta\] \[=\int\limits_{0}^{t}f(t) g(t-\theta) 1 d \theta+\int\limits_{t}^{\infty}f(t) g(t-\theta) 0d \theta\] \[=\int\limits_{0}^{t}f(t) g(t-\theta) d \theta\]

OpenStudy (anonymous):

how can you suppose for 0 to t interval t is greater than theta

OpenStudy (anonymous):

and for t to infnity it is less than theta

OpenStudy (zarkon):

on 0 to t we have \[0<\theta<t\] thus \[t-\theta>0\] and \[u(t-\theta)=1\]

OpenStudy (zarkon):

on t to infinity we have \[t\le \theta\le \infty\] thus \[t-\theta<0\] hence \[u(t-\theta)=0\]

OpenStudy (zarkon):

\[\int\limits_{0}^{t}f(t) g(t-\theta) d \theta\] is a \[d\theta\] integral thus \[0\le \theta\le t\]

OpenStudy (anonymous):

great man !!! realy great ,thank you so much zarkon and myininaya.

OpenStudy (anonymous):

zarkon can you solve one integral, you can see in my questions

myininaya (myininaya):

gj zarkon